### v2

@@ 3.2 ENERGY MEASUREMENTS
Table 3: Comparison of Algorithm Performance and Energy Consumption
Algorithm Segment Length N-Estim. Mean Power [W] Energy [Wh] Time [s] CPU Usage [%] F1 Score
@@ 3.3 EXTENDED USE CASES
3.4 VALIDITY CONSIDERATIONS AND CHALLENGES
@@
In the context of the described approach in diverse domain areas, including fill level detection and sound localization, it is essential to consider potential threats to validity. One challenge relates to the material composition of the bins or objects being monitored. Variations in material properties could affect sensor readings, potentially leading to inaccuracies in classification. Consequently, the SIREC model should be trained to account for these variations. Furthermore, although noise reduction is achieved through spectral subtraction, external factors such as ambient noise or interference from nearby sources might still introduce disturbances, impacting the system’s performance. Additionally, while the experimental setups are designed to simulate real-world conditions, they may not fully capture all the complexities found in realistic environments, possibly limiting the generalizability of the results. Therefore, careful attention to these factors is crucial to ensure the robustness and reliability of our approach across different scenarios and applications.
Future work must also take social and privacy aspects into account. Electronic devices with microphones and the possibility of establishing an internet connection can cause a number of security problems if used incorrectly, some of which we would like to briefly address. A frequently cited problem with insecure IoT scenarios is the emergence of botnets. If our approach were to be used in waste management and thus affect a large number of devices, this possibility of attack must be taken into account. Another attack scenario is unauthorized access to the microphones and the resulting impact on privacy. Attackers can originate from all areas of private and public life: surveillance by governments, use of microphones and data by criminals, e.g. for burglaries, or in the private realm, e.g. for stalking.
4 CONCLUSION
@@
In our study, we have demonstrated the creative potential of AI for addressing environmental challenges. We focused on utilizing sound and a specifically designed ML algorithm to accurately determine the fill levels of containers. Our AI-supported system classifies these levels with over 90% accuracy, highlighting its efficacy in assessing waste container volumes. One of the system’s major strengths is its incorporation of cost-effective components, which significantly enhance scalability and practical application. Furthermore, we explored the energy consumption associated with deploying AI models on resource-constrained devices. Understanding and optimizing energy consumption ensures prolonged battery life, reduced operational costs, and improved sustainability, all of which are vital for edge devices operating in diverse environments with limited power sources. Our approach not only offers a novel method for waste management but also suggests the broader applicability of sound-based classification systems in various scenarios. This methodology could be adapted for other container types or entirely different applications and settings, expanding the potential uses of sound in classification tasks. Looking forward, we hope our findings will inspire further research into the integration of AI technologies within waste management infrastructures. This study opens up new possibilities for how AI can enhance our environmental engagement and resource stewardship, presenting an exciting trajectory for the role of AI in advancing environmental technology.
5 ACKNOWLEDGMENT

### v3

@@ III-B Energy Measurements
TABLE III: Comparison of Algorithm Performance and Energy Consumption
Algorithm Segment Length Ne Mean Power [W] Energy [Wh] Time [s] CPU Usage [%] F1 Score
@@ III-C Extended Use Cases
III-D Validity Considerations and Challenges
@@
In the context of the described approach in diverse domain areas, including fill level detection and sound localization, it is essential to consider potential threats to validity. One challenge relates to the material composition of the bins or objects being monitored. Variations in material properties could affect sensor readings, potentially leading to inaccuracies in classification. Consequently, the SIREC model should be trained to account for these variations. Furthermore, although noise reduction is achieved through spectral subtraction, external factors such as ambient noise or interference from nearby sources might still introduce disturbances, impacting the system’s performance. Additionally, while the experimental setups are designed to simulate real-world conditions, they may not fully capture all the complexities found in realistic environments, possibly limiting the generalizability of the results. Therefore, careful attention to these factors is crucial to ensure the robustness and reliability of our approach across different scenarios and applications. Future work must also take social and privacy aspects into account. Electronic devices with microphones and the possibility of establishing an internet connection can cause a number of security problems if used incorrectly, some of which we would like to briefly address. A frequently cited problem with insecure IoT scenarios is the emergence of botnets. If our approach were to be used in waste management and thus affect a large number of devices, this possibility of attack must be taken into account. Another attack scenario is unauthorized access to the microphones and the resulting impact on privacy. Attackers can originate from all areas of private and public life: surveillance by governments, use of microphones and data by criminals, e.g. for burglaries, or in the private realm, e.g. for stalking.
IV Conclusion
@@
In our study, we have demonstrated the creative potential of AI for addressing environmental challenges. We focused on utilizing sound and a specifically designed ML algorithm to accurately determine the fill levels of containers. Our AI-supported system classifies these levels with over 90% accuracy, highlighting its efficacy in assessing waste container volumes. One of the system’s major strengths is its incorporation of cost-effective components, which significantly enhance scalability and practical application. Furthermore, we explored the energy consumption associated with deploying AI models on resource-constrained devices. Understanding and optimizing energy consumption ensures prolonged battery life, reduced operational costs, and improved sustainability, all of which are vital for edge devices operating in diverse environments with limited power sources. Our approach not only offers a novel method for waste management but also suggests the broader applicability of sound-based classification systems in various scenarios. This methodology could be adapted for other container types, or even applied to entirely different applications and settings, such as surveillance systems for theft protection, thereby broadening the potential applications of sound in classification tasks. Looking forward, we hope our findings will inspire further research into the integration of AI technologies within waste management infrastructures. This study opens up new possibilities for how AI can enhance our environmental engagement and resource stewardship, presenting an exciting trajectory for the role of AI in advancing environmental technology.
V Acknowledgment